OpenStudy (llopiiiiiiiiii):

Factor completely 18x2 − 21x −15.

OpenStudy (andrewmhuff):

Can i get a medal?

OpenStudy (utophii):

\(18 x^2 - 21x - 15 = 3(6x^2 - 7x - 5) = 3(2x+1)(3x-5).\)